OpenStudy (anonymous):

If n is a positive integer, then lim (n-->infinity) (1/n) [1/(1+(1/n)) + 1/(1+(2/n))+...+1/(1+(n/n)] is? the above can be expressed as a.) integral from 0 to 1 of (1/x) dx b.) integral from 1 to 2 of (1/(x+1))dx c.) integral from 1 to 2 of (x)dx d.) integral from 1 to 2 of (2/(x+1)) e.) integral from 1 to 2 of (1/x) please show me how u got the answer thanks a lot!

OpenStudy (e.mccormick):

Been a while since I did calc. This looks a lot like what they use when they go over the Riemann Sum definition of an integral. Is that where you are at in calc?

OpenStudy (amistre64):

might be easier to read in latex

OpenStudy (e.mccormick):

Yah. I was thinking that myself.

OpenStudy (amistre64):

1/n represents (b-a)/n which is the interval from a to b

OpenStudy (amistre64):

they all have intervals of 1, so that doesnt narrow it down

OpenStudy (amistre64):

\[\lim_{n\to inf}~ \sum_{i=1}^{n}\left(\frac{1}{1+\frac in}\right)\frac 1n\] maybe?

OpenStudy (amistre64):

\[Left~side~sums:\lim_{n\to inf}~\sum_{i=0}^{n-1}~f(a+i\frac{b-a}{n})\frac{b-a}{n}\] \[Right~side~sums:\lim_{n\to inf}~\sum_{i=1}^{n}~f(a+i\frac{b-a}{n})\frac{b-a}{n}\]

OpenStudy (amistre64):

i notice some are missing a dx, if that correct, then those can be eliminated for improper format

OpenStudy (amistre64):

c doesnt have division in it, so id scratch that and itd be a toss up between a and b

OpenStudy (amistre64):

when n=0, we have 1/1, and not 1/0 soo im leaning towards b

OpenStudy (amistre64):

when x = a + i/n 1/(x) = 1/(0+i/n) there is also a b version f (b-i(b-a)/n) , for i = 0 to n-1

OpenStudy (amistre64):

1/(1+x) = 1/(1+0+1/n) seems to match best tho

OpenStudy (amistre64):

lol, im sticking with B :) good luck tho